# DreamBench++: A Human-Aligned Benchmark for Personalized Image Generation

### v1

@@ Abstract
Personalized image generation holds great promise in assisting humans in everyday work and life due to its impressive function in creatively generating personalized content. However, current evaluations either are automated but misalign with humans or require human evaluations that are time-consuming and expensive. In this work, we present DreamBench++, a human-aligned benchmark automated by advanced multimodal GPT models. Specifically, we systematically design the prompts to let GPT be both human-aligned and self-aligned, empowered with task reinforcement. Further, we construct a comprehensive dataset comprising diverse images and prompts. By benchmarking 7 modern generative models, we demonstrate that DreamBench++ results in significantly more human-aligned evaluation, helping boost the community with innovative findings.
Figure 1: Overview of DreamBench++. We collect diverse images and prompts, and utilize GPT-4o for automated evaluation aligned with human preference.
@@ 1 Introduction
Driven by the significant advances in large-scale text-to-image (T2I) generative models [85, 82, 8, 83, 69, 88, 116, 14, 27, 21, 22, 7, 49, 24], it is now possible to generate images conditioned on not only arbitrary text prompts but also by given reference images—personalized image generation [86, 28, 55, 115, 53, 29, 6, 17, 48, 15, 111, 98, 109, 67, 44, 105, 66, 103, 16, 99, 126]. In general, to be useful as an artistic creation tool for inspiration or products [113], the following two basic criteria must be fulfilled: i) Prompt following (image & prompt consistency). Generated images must follow the prompt description, which is a requirement shared with vanilla T2I generation [8, 83]. ii) Concept preservation (image & image consistency). For personalized image generation, the concept of the reference image, i.e., the main subject’s semantic details (e.g., facial characters) or high-level abstractions (e.g., overall style), must be preserved in the generated image. For example, a user may want to “imagine his own dog traveling around the world” [86], and the generated dog must be the same as his but traveling.
Figure 2: Qualitative comparison of concept preservation evaluation between DreamBench++ and traditional DINO [13]. DINO often fails to yield human-aligned evaluation while our DreamBench++ succeeds using multimodal GPT models as the evaluator.
To meet the aforementioned requirements, numerous efforts have been devoted. One line of fine-tuning-based works focuses on fine-tuning general T2I models to specialist personalization models by reproducing specific concepts present in training sets [86, 28, 17, 53]. Meanwhile, another line of encoder-based works, instead, achieves concept-preservation by training features adaptation to inject reference image features into a general T2I model [115, 29, 6, 24, 95, 96, 73]. Although remarkable results have been achieved, one question arises: can we comprehensively evaluate these models to figure out which technical route is superior and where to head?
@@
    Human-Aligned  As shown in Fig. 2, traditional metrics like DINO [13] and CLIP [81] often result in significant discrepancies from humans. This is caused by the image similarity measurement nature of DINO and CLIP models, and thus crowd-sourced human evaluation is typically necessary for obtaining a correct quantitative understanding of generated images [54, 52, 112]. Therefore, different from existing works that utilize CLIP and DINO as metrics that may be humanly misaligned, our DreamBench++ demonstrates surprisingly consistent evaluation results aligned with humans. For instance, by evaluating 7 modern models, DreamBench++ achieves 79.64% and 93.18% agreement with human’s evaluation in concept preservation and prompt following capabilities, respectively. Notably, it is +54.1% and +50.7% higher than traditional DINO and CLIP metrics.
@@
    Automated  However, it is non-standardized and expensive to perform high-quality human evaluations, making human evaluations non-trivial to use. To address this challenge, DreamBench++ achieves automated but human-aligned evaluation by using advanced multimodal GPT models as metrics such as GPT-4o [71]. The challenges lie in two aspects: i) prompt design and ii) reasoning procedure for scoring. We systematically standardize the automated GPT evaluation by first designing the evaluation instructions that provide overall task requirements, where language is a general interface for instructing human preference. Inspired by Self-Align [97], we instruct the GPT model to conduct internal thinking that aligns itself for better task and preference understanding. Then, the GPT model provides the summary & planning for the task and scoring criteria, and the final scores are provided with optional chain-of-thought (CoT) [108, 122].
@@
    Diverse  To avoid biased understanding due to the low diversity of evaluation data, DreamBench++ collects a large number of diverse images covering various levels of difficulty categories, from animals and styles that are relatively simple to more challenging human subjects, objects, and non-natural image styles (See Fig. 1). Compared to commonly used DreamBench [86] that only contains 30 subjects and 25 prompts, DreamBench++ scales up the benchmarking data to 150 images and 1,350 prompts, i.e., 5$\times$ and 54$\times$ more than DreamBench. As a result, when evaluating models on this larger dataset, different but more robust conclusions have been obtained that demonstrate an unbiased and more comprehensive evaluation of DreamBench++.
Takeaways  Here, we offer some insightful findings that we obtain by evaluating 7 modern personalized T2I models: i) DINO-based rating puts tremendous weight on resemblance in overall shape or color and neglects detailed visual features, which makes it a sub-optimal metric for personalized image generation evaluation; ii) The main goal of personalized image generation is to find the Pareto optimal solution that balances concept preservation and prompt following. Among these 7 models, DreamBooth [86] is the best in overall performance, preserving highly detailed visual features while keeping close adherence with text prompt; iii) Currently, personalized T2I models perform well on animal and style category. However, they have not performed well in the human category due to sensitivity to facial details and in the object category due to its diversity. Existing works [105, 100, 114, 115, 111] focuses on preserving facial features to address the former problem, but the latter has not been extensively explored.
@@ 2 DreamBench++
We introduce DreamBench++, a human-aligned, automated, and diverse benchmark that evaluates the two capabilities of personalized image generation models. In the following, we describe how we construct DreamBench++ from two aspects: prompts and data.
Figure 3: Overall procedure of prompting GPT-4o for automated evaluation. The evaluation instructions are meta-prompting information written by humans, including task description, scoring criteria, scoring range, and format specification. Then, GPT-4o is prompted with reasoning instructions to perform internal thinking that provides a self-aligned task summary and planning. Finally, all prompts and reasoning outputs are joined with image samples for score outputs.
2.1 Prompting Multimodal GPT for Automated & Human-Aligned Benchmarking
@@
It is challenging to obtain a solid quantitative understanding of generated models, especially when evaluating visual contents like images that rely on human evaluations [54, 52]. Thus, it is critical to achieve automated evaluation by utilizing multimodal GPT models, which are trained particularly in the principle of aligning with human preference [72, 20, 71, 70]. This is evidenced by the recent progress achieved by Wu et al., which demonstrates that GPT-4V [70] can serve as a human-aligned text-to-3D generation evaluator. However, as pointed out by Zhang et al. and Ku et al., multimodal GPT models often fall short in evaluating personalized image generation—often more challenging when distinguishing subtle difference for concept preservation assessment using GPT—still underexplored. To tackle this issue, we detail how we systematically design the prompt of multimodal GPT (GPT-4o [71], by default) for human alignment reinforcement but also improve the reasoning progress that helps the GPT models to be more self-aligned, introduced as follows.
@@
Compare or rate?  There are typically two schemes for quantitatively evaluating generative models in human evaluations: rating and comparison [121, 124]. The rating scheme requires human reviewers to assign an absolute score to each instance, while the comparison scheme asks human reviewers to express a relative preference among different instances. Though effective as the comparison scheme is when humans are involved, we find that there are two critical issues. i) Positional Bias: the scoring results of GPT-4V/GPT-4o is sensitive to the order in which images are presented [70, 104, 107, 121, 110, 124], making it unsuitable for comparison scheme. ii) Quadratic Complexity: As the number of methods increases, the number of essential evaluation runs for numerical rating increases linearly, while the number of comparative assessments increases quadratically. Therefore, direct numerical rating is more efficient and scalable when evaluating multiple methods. Hence, in this work, we adhere to the rating scheme, and we establish a 5-level rating scheme where scores are integers ranging from 0 (very poor) to 4 (excellent).
@@
Figure 4: Dataset construction process of our DreamBench++. We start by obtaining keywords through GPT generation, existing datasets, and human proposals. Next, we collect corresponding images from the internet. These images are then filtered to remove low-quality ones through both model and human assessment. The remaining high-quality images are used as input for GPT-4o to generate text prompts of varying difficulty levels.
@@
Reasoning Instructions  Given the evaluation instructions, it is crucial to reinforce the alignment with both the human instruction and itself to largely leverage the pretrained knowledge. To this end, we adopt a 2-step evaluation policy as follows: i) Internal Thinking: Inspired by Self-Align [97], we introduce internal thinking to strengthen task understanding and instruction following capabilities. Specifically, we prompt the GPT model by asking if it understands the task or not and let it summarize the task. ii) Summary & Planning: According to the given internal thinking instruction, the GPT will summarize and plan for the evaluation task itself. It can also be viewed as a generalized form of chain-of-thought reasoning [108, 122]. The complete procedure is illustrated in Fig. 3.
@@ 2.2 Scaling Up Personalized Image Generation Benchmarking
Pioneering works like DreamBooth [86] and SuTI [17] have successfully set up baseline datasets for the evaluation of personalized image generation, and DreamBench++ follows them to categorize images into three types: ❶ objects, ❷ living subjects, and ❸ styles. However, due to the small-scale nature of DreamBench, it is limited as some methods may converge well on its samples while performing unsatisfactorily on other data. To avoid this possible biased evaluation, we scale up the benchmarking data by increasing both image numbers and diversity.
Figure 5: Data distribution visualization. (a) Images comparison between DreamBench and DreamBench++ using t-SNE [101, 77]. (b) Image and prompt distribution of DreamBench++.
Data Construction from Internet  There are broad images on the Internet, and many datasets are constructed from it [90, 47]. DreamBench++ mainly collects images from Unsplash [3], Rawpixel [2], and Google Image Search [1], along with contributions from individuals with authorized permissions. Each image’s copyright status has been verified for academic suitability. As shown in Fig. 4, we collect and construct high-quality data in DreamBench++ by following 3 steps:
@@
  • Internet Images Collection  Given selected keywords, we retrieved corresponding images from Unsplash, Rawpixel, and Google Image Search. To filter out images unsuitable for personalized image generation, SAM [50] is applied to identify subject regions in images and discard those with too small subject areas. Human annotators will then filter out images with noisy backgrounds. Curated images were cropped to centralize the subject, resulting in two images per keyword. Keywords that fail to yield suitable images will be discarded in this process.
  • Prompt Generation  After image collection, 9 text prompts per image were generated using GPT-4o, designed to cover a range of difficulties: 4 prompts for ❶ photorealistic styles, 3 for ❷ non-photorealistic styles, and 2 for ❸ complicated & imaginative contents. To align with established evaluation methods, we use few-shot prompts selected from PartiPrompts [116]. Human calibration ensures that all generated prompts are ethical and without flaws. As a result, the construction process finally yields 150 high-quality images and 1,350 prompts.
Diversity Visualization  When collecting images, we notice that Internet images possess a bias towards photorealistic styles. To diversify, various non-photorealistic styles are enlisted, and human annotators are tasked to gather images for each style, including anime, sketches, traditional Chinese paintings, artworks, and cartoon characters from games. Then, a manual selection process ensures a balanced distribution of images across subject classes and between photorealistic and non-photorealistic styles. In Fig. 5(a), we visualize the t-SNE [101, 77] of images from DreamBench and DreamBench++, which demonstrates the superiority of DreamBench++ in diversity. Besides, Fig. 5(b) presents the detailed statistics of the image distribution in DreamBench++.
@@ 3 Experiments
In this section, we provide evidence that DreamBench++ can more accurately reflect the capabilities of existing personalized image generation methods with its diverse and comprehensive data set. Additionally, we demonstrate that our systematic prompts-based GPT is a superior evaluator compared to existing metrics by analyzing their degree of alignment with human ratings. In order to further facilitate benchmarking research with GPT models, we conduct a comprehensive ablation study to show the designing key points for more human-aligned automated GPT evaluation.

### v2

@@ Abstract
Personalized image generation holds great promise in assisting humans in everyday work and life due to its impressive ability to creatively generate personalized content across various contexts. However, current evaluations either are automated but misalign with humans or require human evaluations that are time-consuming and expensive. In this work, we present DreamBench++, a human-aligned benchmark that advanced multimodal GPT models automate. Specifically, we systematically design the prompts to let GPT be both human-aligned and self-aligned, empowered with task reinforcement. Further, we construct a comprehensive dataset comprising diverse images and prompts. By benchmarking 7 modern generative models, we demonstrate that DreamBench++ results in significantly more human-aligned evaluation, helping boost the community with innovative findings.
Figure 1: Overview of DreamBench++. We collect diverse images and prompts as the base dataset and utilize GPT-4o for automated evaluation aligned with human preference. To evaluate personalized image generation, DreamBench++ mainly focuses on image & image concept preservation and text & image prompt following assessments.
@@ 1 Introduction
Driven by the significant advances in large-scale text-to-image (T2I) generative models (Rombach et al., 2022; Ramesh et al., 2021; Betker et al., 2023; Ramesh et al., 2022; Nichol et al., 2022; Saharia et al., 2022b; Yu et al., 2022; Chang et al., 2023; Gafni et al., 2022; Ding et al., 2021; 2022; Balaji et al., 2022; Kang et al., 2023; Dong et al., 2024), it is now possible to generate images conditioned on not only arbitrary text prompts but also by given reference images—personalized image generation (Ruiz et al., 2023; Gal et al., 2023a; Li et al., 2023a; Ye et al., 2023; Kumari et al., 2023; Gal et al., 2023b; Arar et al., 2023; Chen et al., 2023c; Jia et al., 2023; Chen et al., 2023a; Xiao et al., 2023; Tewel et al., 2023; Wei et al., 2023; Ma et al., 2023; Hua et al., 2023; Wang et al., 2024b; Lv et al., 2024; Wang et al., 2024a; Chen et al., 2023b; Tumanyan et al., 2023; Zhou et al., 2024; Tan et al., 2024; He et al., 2024c; Wang et al., 2024c; Wu et al., 2024a; He et al., 2024a; Xiao et al., 2024; Arar et al., 2024; Huang et al., 2024b; a; Pang et al., 2024a; b; Qiu et al., 2023; Hu et al., 2024). In general, to be useful as an artistic creation tool for inspiration or products (Yacoubian, 2022; Xie et al., 2024), the following two basic criteria must be fulfilled: i) Prompt following (image & prompt consistency). Generated images must follow the prompt description, which is a requirement shared with vanilla T2I generation (Betker et al., 2023; Ramesh et al., 2022). ii) Concept preservation (image & image consistency). For personalized image generation, the concept of the reference image, i.e., the main subject’s semantic details (e.g., facial characters) or high-level abstractions (e.g., overall style), must be preserved. For example, a user may want to “imagine his own dog traveling around the world” (Ruiz et al., 2023), and the generated dog must be the same as his but traveling.
Figure 2: Qualitative comparison of concept preservation evaluation between DreamBench++ and traditional DINO (Caron et al., 2021). DINO often fails to yield human-aligned evaluation while our DreamBench++ succeeds using multimodal GPT models as the evaluator.
To meet the aforementioned requirements, numerous efforts have been devoted. One line of fine-tuning-based works focuses on fine-tuning general T2I models to specialist personalization models by reproducing specific concepts present in training sets (Ruiz et al., 2023; Gal et al., 2023a; Chen et al., 2023c; Kumari et al., 2023). Meanwhile, another line of encoder-based works, instead, achieves concept-preservation by training features adaptation to inject reference image features into a general T2I model (Ye et al., 2023; Gal et al., 2023b; Arar et al., 2023; Dong et al., 2024; Sun et al., 2024a; b; Pan et al., 2024). Despite remarkable progress, one question arises: can we comprehensively evaluate these models to figure out which technical route is superior and where to head?
Figure 3: Overall procedure of prompting GPT-4o for automated evaluation. Human-written meta-prompts provide task description, scoring criteria, range, and format. GPT-4o then follows reasoning instructions for self-aligned task summarization and planning. Finally, all prompts and reasoning outputs are combined with image samples for score generation.
@@
    Human-Aligned  As shown in Fig. 2, traditional metrics like DINO (Caron et al., 2021) and CLIP (Radford et al., 2021) often result in significant discrepancies from humans. This is caused by the image similarity measurement nature of DINO and CLIP models, and thus crowd-sourced human evaluation is typically necessary for obtaining a correct quantitative understanding of generated images (Lee et al., 2023; Ku et al., 2024; Xu et al., 2023). Therefore, different from existing works that utilize CLIP and DINO as metrics that may be humanly misaligned, our DreamBench++ demonstrates surprisingly consistent evaluation results aligned with humans. For instance, by evaluating 7 modern models, DreamBench++ achieves 79.64% and 93.18% agreement with human’s evaluation in concept preservation and prompt following capabilities, respectively. Notably, it is +32.59% and +37.23% higher than traditional DINO and CLIP metrics.
@@
    Automated  However, it is non-standardized and expensive to perform high-quality human evaluations. To address this challenge, DreamBench++ achieves automated but human-aligned evaluation by using advanced multimodal GPT models such as GPT-4o (OpenAI, 2024) as metrics. The challenges lie in two aspects: i) prompt design and ii) reasoning procedure for scoring. We systematically standardize the automated GPT evaluation by first designing the evaluation instructions that provide overall task requirements, where language is a general interface for instructing human preference. Inspired by Self-Align (Sun et al., 2023), we instruct GPT to conduct internal thinking that aligns itself for better task and preference understanding. Then, GPT provides the summary & planning for the task and scoring criteria, and the final scores are provided with optional chain-of-thought (CoT) (Wei et al., 2022; Zhang et al., 2023d).
@@
    Diverse  To avoid bias from low-diversity evaluation data, DreamBench++ compiles a wide range of images, covering varying levels of difficulty from simpler animals and styles to more complex human subjects, objects, and non-natural styles (see Fig. 1). Unlike DreamBench (Ruiz et al., 2023), which includes only 30 subjects and 25 prompts, DreamBench++ significantly expands the dataset to 150 images and 1,350 prompts—5$\times$ and 54$\times$ more, respectively. While CustomConcept101 (Kumari et al., 2023) offers 101 subjects, its diversity is limited by repetitive image categories and a focus on photorealistic styles, with simple prompts that restrict its ability to evaluate models on more complex tasks. Consequently, DreamBench++ enables more robust and comprehensive conclusions in model evaluation.
Takeaways  We present some insightful findings from evaluating seven modern personalized T2I models: i) DINO-based ratings prioritize overall shape and color over detailed features, making them suboptimal for evaluating personalized image generation; ii) The primary goal is to achieve a Pareto optimal balance between concept preservation and prompt adherence. Among the models, DreamBooth (Ruiz et al., 2023) excels in preserving detailed visual features while closely following text prompts; iii) Current models perform well in animal and style categories but struggle with human images due to sensitivity to facial details and diverse object categories. While existing work (Wang et al., 2024b; Valevski et al., 2023; Yan et al., 2023; Ye et al., 2023; Xiao et al., 2023) addresses facial feature preservation, the challenge of object diversity remains underexplored.
@@ 2 DreamBench++
2.1 Prompting GPT for Automated & Human-Aligned Benchmarking
@@
It is challenging to obtain a solid quantitative understanding of generated models, especially when evaluating visual contents that rely on human evaluations (Lee et al., 2023; Ku et al., 2024). Thus, it is critical to achieve automated evaluation by utilizing multimodal GPT models, which are trained particularly in the principle of aligning with human preference (Ouyang et al., 2022; Christiano et al., 2017; OpenAI, 2024; 2023). This is evidenced by the recent progress achieved by Wu et al., which demonstrates that GPT-4V (OpenAI, 2023) can serve as a human-aligned text-to-3D generation evaluator. However, as pointed out by Zhang et al. and Ku et al., multimodal GPT models often fall short in evaluating personalized image generation—often more challenging when distinguishing subtle difference for concept preservation assessment using GPT—still underexplored. To tackle this issue, we detail how we systematically design the prompt of multimodal GPT (GPT-4o (OpenAI, 2024), by default) for human alignment reinforcement but also improve the reasoning progress that helps the GPT models to be more self-aligned, introduced as follows.
@@
Compare or rate?  There are typically two schemes for quantitatively evaluating generative models in human evaluations: rating and comparison (Zhang et al., 2023c; Zheng et al., 2023). The rating scheme requires human reviewers to assign an absolute score to each instance, while the comparison scheme asks human reviewers to express a relative preference among different instances. Though effective as the comparison scheme is when humans are involved, we find that there are two critical issues. i) Positional Bias: the scoring results of GPT-4V/GPT-4o is sensitive to the order in which images are presented (OpenAI, 2023; Wang et al., 2023a; b; Zhang et al., 2023c; Wu et al., 2024b; Zheng et al., 2023), making it unsuitable for comparison scheme. ii) Quadratic Complexity: As the number of methods increases, the number of essential evaluation runs for numerical rating increases linearly, while the number of comparative assessments increases quadratically. Therefore, direct numerical rating is more efficient and scalable when evaluating multiple methods. Hence, in this work, we adhere to the rating scheme, and we establish a 5-level rating scheme where scores are integers ranging from 0 (very poor) to 4 (excellent).
@@
Figure 4: Dataset construction process of our DreamBench++. We generate keywords using GPT-4, existing datasets, and human input, then collect corresponding web images. After filtering out low-quality images via model and human assessment, the high-quality ones are used as input for GPT-4o to generate text prompts of varying difficulty.
@@
Reasoning Instructions  Given the evaluation instructions, it is crucial to reinforce the alignment with both the human instruction and itself to largely leverage the pretrained knowledge. To this end, we adopt a 2-step evaluation policy: i) Internal Thinking: Inspired by Self-Align (Sun et al., 2023), we introduce internal thinking to strengthen task understanding and instruction following capabilities. Specifically, we prompt the GPT model by asking if it understands the task or not and let it summarize the task. ii) Summary & Planning: According to the given internal thinking instruction, the GPT will summarize and plan for the evaluation task itself. It can also be viewed as a generalized form of chain-of-thought reasoning (Wei et al., 2022; Zhang et al., 2023d). See Fig. 3.
@@ 2.2 Scaling Up Personalized Image Generation Benchmarking
Pioneering works like DreamBooth (Ruiz et al., 2023), SuTI (Chen et al., 2023c) and CustomConcept101 (Kumari et al., 2023) have successfully set up baseline datasets evaluating personalized image generation, and DreamBench++ follows them to categorize images into three types: ❶ objects, ❷ living subjects, and ❸ styles. However, DreamBench’s small scale and CustomConcept101’s limited diversity may lead to biased evaluations, as some methods may converge well on its samples while performing unsatisfactorily on other data. To mitigate this, we scale up data by increasing both the number and diversity of images.
Figure 5: Data distribution visualization. (a) Images comparison between DreamBench and DreamBench++ using t-SNE. (b) Image and prompt distribution of DreamBench++.
Data Construction from Internet  The internet offers vast images, with many datasets built from it (Schuhmann et al., 2021; Jia et al., 2021). DreamBench++ primarily sources images from Unsplash (uns, ), Rawpixel (raw, ), and Google Image Search (goo, ), along with authorized individual contributions. Each image’s copyright status has been verified for academic suitability. As shown in Fig. 4, we collect and curate high-quality data to ensure robustness and diversity.
@@
  • Internet Images Collection  Given selected keywords, we retrieved images from Unsplash, Rawpixel, and Google Image Search. To filter out unsuitable images, SAM (Kirillov et al., 2023) identifies subject regions, discarding those with small subject areas. Human annotators remove images with noisy backgrounds. Curated images are then cropped to centralize the subject, yielding two images per keyword. Keywords without suitable images are discarded.
  • Prompt Generation  After image collection, 9 text prompts per image were generated using GPT-4o, designed to cover a range of difficulties: 4 prompts for ❶ photorealistic styles, 3 for ❷ non-photorealistic styles, and 2 for ❸ complicated & imaginative contents. To align with established evaluation methods, we use few-shot prompts selected from PartiPrompts (Yu et al., 2022). Human calibration ensures that all generated prompts are ethical and without flaws. As a result, the construction process finally yielded 150 high-quality images and 1,350 prompts.
Diversity Visualization  Internet images are numerous. However, there is a bias towards photorealistic styles. To diversify, various non-photorealistic styles are enlisted, and human annotators are tasked to gather images for each style, including anime, sketches, traditional Chinese paintings, artworks, and cartoon characters from games. Then, a manual selection process ensures a balanced distribution across subject classes and between photorealistic and non-photorealistic styles. In Fig. 5(a), we visualize the t-SNE (Van der Maaten & Hinton, 2008; Poličar et al., 2019) of images from DreamBench and DreamBench++, which demonstrates the superiority of DreamBench++ in diversity. Besides, Fig. 5(b) presents the detailed image distribution in DreamBench++.